OpenStudy (anonymous):

When mass M is at the position shown, it is sliding down the inclined part of a slide at a speed of 1.95 m/s. The mass stops a distance S2 = 2.15 m along the level part of the slide. The distance S1 = 1.10 m and the angle theta = 37.1 degrees. Calculate the coefficient of kinetic friction for the mass on the surface.

OpenStudy (anonymous):

Find the deceleration from friction: V2^2 = V1^2 - 2ad Then add up force of friction and the component of mg along the plane and set it equal to ma.

OpenStudy (anonymous):

i dont understand.. can you show me the steps?

OpenStudy (anonymous):

After you find the deceleration from the equation given, add up the forces and set it equal to ma: Frictional force: \[F_{f} = \mu_{k}mg\cos{\theta}\] The component of mg in the direction of the plane = mgSin(theta) Draw a picture of all these forces, including the Normal force which gives the equation for friction.

OpenStudy (anonymous):

Can you figure it out from here?

OpenStudy (anonymous):

Sorry, the first equation should be: V2^2 = V1^2 + 2ad That will give you a negative acceleration. Then look at the direction of the frictional force and the component of mg along the incline to determine if they should be positive or negative.